Bubba’s Beers has won the beer concession for Coconut Grove Fair. This means that Bubba can locate one beer stand in the middle of a one mile street, where fair takes place. The stand costs $500 to build and staff. He expects
N = 3000 people per day, each of whom wants to buy one beer, and has a maximum willingness to pay of $3. Assume that, when the thirst hits, consumers are evenly spread out along the street. In addition to the price, consumers also
experience disutility from having to walk to get the beer, at the (round-trip) rate of $2 per mile of distance to the nearest stand. Suppose that Bubba’s marginal costs are constant and equal to $1.5 per beer.
(a) At x miles to the left of the beer staff, the furthest consumer (to the left) to buy beer from Bubba is located. Write
the demand function for Bubba’s beer in terms x. Q(P) = _
(b) Find the price charges for a beer, and, location of consumer x, and profits. How many people
are served on the street? P* = _ πL
* = _ Q* =
1
Expert's answer
2019-10-17T10:16:09-0400

(a) The demand function for Bubba’s beer in terms x. Q(P) = 3×3000 + 2×3000x - P = 9000 + 6000x - P.

(b) TC = 500 + 1.5Q,

MC = TC'(Q) = 1.5,

MR = TR'(Q) = 9000 + 6000x - 2Q.

MR = MC when optimal output is produced.

9000 + 6000x - 2Q = 1.5,

Q = 4500 + 3000x - 0.75,

P = 9000 + 6000x - (4500 + 3000x - 0.75) = 4500.75 - 3000x.

TP = TR - TC.